OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. A. y = 4x - 11 B. y = 4x + 11 C. y = 4x + 7 D. y = 4x - 7

OpenStudy (anonymous):

B

hero (hero):

At least provide steps. Don't just give the answer.

hero (hero):

Insert given values into y = mx + b: -1 = 4(-3) + b Now solve for the unknown variable, b: -1 = -12 + b 12 - 1 = b 11 = b Therefore the equation of the line is y = 4x + 11

OpenStudy (anonymous):

thanks (: